Show that R = {a + b√-3 a, b belongs to Z} i s not a
unique factorization domain by expressing
4 as a product of two irreducible elements in
R in two different ways.
1
Expert's answer
2020-12-02T01:43:33-0500

We can clearly see that

"4 = 2\cdot 2 = (1+1\sqrt{-3})\cdot (1-1\sqrt{-3})"

First let's prove that they are irreducible.

We can view elements of "R" as complex numbers (obviously) and thus we know that if "2=(a+b\sqrt{-3})(c+d\sqrt{-3})" we have "(a^2+3b^2)(c^2+3d^2)=4" and as the last expression is an equation in "\mathbb{Z}" we can easily conclude that either "a^2+3b^2=1," or "c^2+3d^2=1" and thus we have "2= (\pm2) \cdot (\pm1)" and "2" is irreducible.

Now let's prove that "1\pm\sqrt{-3}" are irreducible. By the same argument of absolute values in "\mathbb{C}" we find that "1+\sqrt{-3} = (\pm1)\cdot (\pm1 \cdot (1+\sqrt{-3}))" and same for "1-\sqrt{-3}" . Therefore they are also irreducible.

In addition, while proving this we have studied all unit elements : "\pm1" and so "1+\sqrt{-3}" and "2" are not associated. Thus we conclude that "R" is not a unique factorization domain.